# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

### v4

@@ Appendix A Simulation Plots
This appendix compiles all simulation plots of $\Delta\gamma$ as functions of $v_{2}$ analyzed by the ESS and ESE methods from physics models as well as from toy models.
@@
Figure A.2: avfd ESE results. The same avfd data as in Fig. A.2 is used. The $\Delta\gamma$ is plotted as a function of $mean{v_{2}}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<2$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). The red line is a first-order polynomial fit to all data points.

### v5

@@ Appendix A Simulation Plots
This appendix compiles, in Figs. A.2-A.14, all simulation plots of $\Delta\gamma$ as functions of $v_{2}$ analyzed by the ESS and ESE methods from physics models as well as from toy models.
@@
Figure A.2: avfd ESE results. The same avfd data as in Fig. A.2 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<2$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). The red line is a first-order polynomial fit to all data points.